y"+2y'+y=3e^-x√x+1 find the general solution
1
Expert's answer
2020-10-27T18:38:06-0400

"y'' + 2y' + y = 3{e^{ - x}}\sqrt {x + 1}"

Let's find "{y_c}" , using auxiliary equation:

"\begin{array}{l}
{k^2} + 2k + 1 = 0\\
{(k + 1)^2} = 0\\
{k_1} = {k_2} = - 1
\end{array}"

Therefore "{y_c} = {C_1}{e^{ - x}} + x{C_2}{e^{ - x}}"

Let

"{C_1} = f(x),\,{C_2} = g(x)"

Then

"y = f{e^{ - x}} + xg{e^{ - x}} \Rightarrow y' = f'{e^{ - x}} - f{e^{ - x}} + g({e^{ - x}} - x{e^{ - x}}) + g'x{e^{ - x}}"

"\begin{array}{l}
Let\\
f'{e^{ - x}} + g'x{e^{ - x}} = 0\,\,\,(*)
\end{array}"

Then

"y' = - f{e^{ - x}} + g({e^{ - x}} - x{e^{ - x}}) \Rightarrow y'' = - f'{e^{ - x}} + f{e^{ - x}} + g'({e^{ - x}} - x{e^{ - x}}) + g( - 2{e^{ - x}} + x{e^{ - x}})"

Substituting into our equation for y, y', y'', we have:

"\begin{array}{l}
 - f'{e^{ - x}} + f{e^{ - x}} + g'({e^{ - x}} - x{e^{ - x}}) + g( - 2{e^{ - x}} + x{e^{ - x}}) + 2\left( { - f{e^{ - x}} + g({e^{ - x}} - x{e^{ - x}})} \right) + \\
 + f{e^{ - x}} + xg{e^{ - x}} = 3{e^{ - x}}\sqrt {x + 1} 
\end{array}" "- f'{e^{ - x}} + g'({e^{ - x}} - x{e^{ - x}}) = 3{e^{ - x}}\sqrt {x + 1}"

with(*) we have

"\left\{ \begin{array}{l}
 - f'{e^{ - x}} + g'({e^{ - x}} - x{e^{ - x}}) = 3{e^{ - x}}\sqrt {x + 1} \\
f'{e^{ - x}} + g'x{e^{ - x}} = 0\,
\end{array} \right."

From 2-nd equation:

"f' = - g'x"

Then

"g'x + g' - g'x = 3\sqrt {x + 1} \Rightarrow g' = 3\sqrt {x + 1} \Rightarrow g = 2\sqrt {x + 1} (x + 1) + C"

"\begin{array}{l}
f' = - g'x = - 3x\sqrt {x + 1} \Rightarrow f = - \frac{2}{5}\left( {3x - 2} \right)\left( {x + 1} \right)\sqrt {x + 1} + {C_2}\\
y = f{e^{ - x}} + xg{e^{ - x}} = \left( { - \frac{2}{5}\left( {3x - 2} \right)\left( {x + 1} \right)\sqrt {x + 1} + {C_2}} \right){e^{ - x}} + x\left( {2\sqrt {x + 1} (x + 1) + C} \right){e^{ - x}}
\end{array}"